# Expansion-sensitive coupling of a local quantum system in de Sitter cosmology

### v1

@@ I Introduction
Let us briefly develop the idea and introduce the concept we refer to as dynamics. For the sake of the example, consider a scalar field $\phi$ in a globally hyperbolic, spatially homogeneous and isotropic spacetime. We begin by describing the system at time equal $t_{0}$ by the corresponding Fourier coefficients and Fourier modes and promote these coefficients to operators living in Fock space, or Hilbert space in the case of one particle. Note, however, that if we choose to describe the system at time $t\neq t_{0}$ instead, different (or not necessarily the same) Fourier coefficients are obtained. We can choose to work with the same Fock space as previously and note that both field descriptions are related by a Bogoliubov transformation. Hence, by choosing the same representation for both descriptions of the field, this latter transformation is found by solving the corresponding equations of motion, assuming the initial conditions of the system are well-posed. We can thus understand this Bogoliubov transformation as dynamics. Moreover, an important observation is to be made. Indeed, to simplify the equations of motion, in cosmological spacetimes it is convenient to introduce what is known as the conformal field $\chi$, equal (or proportional) to the scalar field times the scale factor, and even though the ground states of the two fields correspond to the same vacuum, their respective dynamics are different. This fact usually does not pose any concern, since the dynamics of the two fields are closely related; i.e. by knowing the solution $\chi$ at time equal $t_{0}$, we also know $\phi$ at time equal $t_{0}$. Nevertheless, it is noteworthy that, in the case of a massless field, the action of the conformal field can be recast as that of a field in flat spacetime, thus decoupled from geometrical dynamical contributions. Similarly, in the case of a massive scalar field, we arrive at the same results for the kinetic term in the Lagrangian of the conformal field, alongside acquiring an effective mass that depends on the scale factor. In fact, one of the primary motivations of introducing the conformal field is to be able to identify a canonical kinetic [^1] We refer to a canonical term $X$ for a scalar field $\phi$ as the term defined by $\tfrac{1}{2}\bigl{(}\dot{\phi}^{2}-(\partial_{i}\phi)^{2}\bigr{)}$, where the overdot denotes a time derivative and $\partial_{i}$ the space derivatives. The canonical scalar field corresponds then to the choice of a densitized Lagrangian $\mathcal{L}=X-V(\phi)$, where $V(\phi)$ is the potential of the field [14]. term of the theory [15]. Therefore, one could think of the dynamics of the conformal field, governed by the corresponding canonical kinetic term, as the proper dynamics of the field.
@@
Turning back our discussion to the context of detectors, given the rising interest in time-dependent couplings for detectors, it is of high interest to consider the coupling to the conformal field, since mathematically it is equivalent to considering the usual coupling to the field $\phi$ with a background-dependent and hence time-dependent scaling in that coupling. In particular, we shall study the different characteristics of the particle measuring in de Sitter spacetime resulting from these fields with distinct dynamics —one with standard dynamics and the other with proper dynamics— and determine which corresponding detector reading better aligns with our intuitive notion of a particle. Moreover, we shall require the time dependence of the detector coupling to be in a functional (and thus direct) form with respect to the spacetime, to allow that it can be fully specified by endowing the detector with the capacity to perform a direct reading of the spacetime without any knowledge in advance of its behaviour. In that way, one could think of processing the outcome signal of a more classical or conventional detector. Hence, it is worth highlighting that in this sense, through these independent readings of spacetime, the new detector under consideration will be capable of absorbing the possible excitations caused by the background dynamics and be well-adapted to the expanding cosmology.
@@ Acknowledgments
A.T.-C. is grateful to Antonio Vicente-Becerril and Bruno Falcone for constructive discussions on numerics and to the Gravity Laboratory Group of the University of Nottingham for the interesting discussions on the subject. G.A.M.M. acknowledges financial support from Project No. PID2020-118159GB-C41 and Project No. PID2023-149018NB-C41, both funded by MCIN/AEI/10.13039/501100011033/, and Project No. 2024AEP005 from the Spanish National Research Council (CSIC). The work of J.L. was supported by the United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.

### v2

@@ I Introduction
Let us briefly develop the idea and introduce the concept we refer to as dynamics. For the sake of the example, consider a scalar field $\phi$ in a globally hyperbolic, spatially homogeneous and isotropic spacetime. We begin by describing the system at time equal $t_{0}$ by the corresponding Fourier coefficients and Fourier modes and promote these coefficients to operators living in Fock space, or Hilbert space in the case of one particle. Note, however, that if we choose to describe the system at time $t\neq t_{0}$ instead, different (or not necessarily the same) Fourier coefficients are obtained. We can choose to work with the same Fock space as previously and note that both field descriptions are related by a Bogoliubov transformation. Hence, by choosing the same representation for both descriptions of the field, this latter transformation is found by solving the corresponding equations of motion, assuming the initial conditions of the system are well-posed. We can thus understand this Bogoliubov transformation as dynamics. Moreover, an important observation is to be made. Indeed, to simplify the equations of motion, in cosmological spacetimes it is convenient to introduce what is known as the conformal field $\chi$, equal (or proportional) to the scalar field times the scale factor, and even though the ground states of the two fields correspond to the same vacuum, their respective dynamics are different. This fact usually does not pose any concern, since the dynamics of the two fields are closely related; i.e. by knowing the solution $\chi$ at time equal $t_{0}$, we also know $\phi$ at time equal $t_{0}$. Nevertheless, it is noteworthy that, in the case of a massless field, the action of the conformal field can be recast as that of a field in flat spacetime, thus decoupled from geometrical dynamical contributions. Similarly, in the case of a massive scalar field, we arrive at the same results for the kinetic term in the Lagrangian of the conformal field, alongside acquiring an effective mass that depends on the scale factor. In fact, one of the primary motivations of introducing the conformal field is to be able to identify a canonical kinetic [^1] We refer to a canonical term $X$ for a scalar field $\phi$ as the term defined by $\tfrac{1}{2}\bigl(\dot{\phi}^{2}-(\partial_{i}\phi)^{2}\bigr)$, where the overdot denotes a time derivative and $\partial_{i}$ the space derivatives. The canonical scalar field corresponds then to the choice of a densitized Lagrangian $\mathcal{L}=X-V(\phi)$, where $V(\phi)$ is the potential of the field [16]. term of the theory [17]. Therefore, one could think of the dynamics of the conformal field, governed by the corresponding canonical kinetic term, as the proper dynamics of the field [^2] For an argument based on unitary implementable dynamics, see Ref. [18]..
The study of quantum field theory in curved spacetimes, such as in cosmological backgrounds, covers a wide range of areas, some of which have in fact been very actively explored recently. One active direction concerns the characterisation and harvesting of entanglement in quantum fields [19, 20, 21], with some works exploring time-dependent detector couplings as a tool to manage causal separation and finite-time interactions, e.g. Refs. [22, 23, 24, 25, 26, 27, 28, 29, 30]. Another notable line of work focuses on the unitary evolution of quantum states in dynamical geometries. In particular, several works argue that the unitary implementability of field dynamics provides a criterion for selecting a unique quantum representation (up to unitary equivalence) [31, 18, 32, 33] and argue that it is important for defining a consistent relationship between the Schrödinger and Heisenberg pictures in cosmological and black-hole settings [34]. These advances provide a strong motivation for the present work, as both perspectives —time-dependent couplings and unitary implementable field dynamics— shed light on different aspects of quantum field evolution in dynamical backgrounds, and may prove relevant for pre-inflationary analyses such as that presented in Ref. [35].
@@
Therefore, turning back to the context of detectors, and given the growing interest in this field, it becomes natural to investigate the effects of introducing a time-dependent scaling in the detector’s coupling. In previous approaches in this area, this time dependence has been introduced by hand, e.g. to model detectors sensitive to the Berry phase rather than excitation responses [36], to recover the long-interaction limit [37, 38], or to account for finite-time interactions motivated by physical considerations such as distinguishing early- and late-time behaviour [39, 40, 41]. In contrast, following our previous discussion, we will consider here a background-dependent coupling, and in particular, the coupling to the conformal field, since mathematically it is equivalent to considering the usual coupling to the field $\phi$ with a background-dependent and hence time-dependent scaling in that coupling. More specifically, we will study the different characteristics of the particle measuring in de Sitter spacetime resulting from these two fields with distinct dynamics —one with standard dynamics and the other with proper dynamics. In this way, we will determine which corresponding detector reading better aligns with our intuitive notion of a particle. Moreover, our requirement that the time dependence of the detector coupling be in a direct functional form with respect to the spacetime allows that it can be fully specified by endowing the detector with the capacity to perform a direct reading of the background spacetime without any knowledge in advance of its behaviour. In essence, for a time dependence of the coupling restricted to come entirely from the dynamical background, we could think of a detector device able to measure this background simultaneously and independently, and then engineer or design a data processing system that incorporates this information into the final signal by means of a Fourier convolution. This convolution would have the same effect as if the original detector’s response function were multiplied by the background dependent function of our choice. In total, the set-up would modify the detector’s outcome in terms of the background signal without a prior knowledge of how this background evolves or the use of any externally prescribed control profile required to emulate its evolution. Hence, in this sense, through these independent readings of spacetime, the new detector under consideration will be capable of absorbing the possible excitations caused by the background dynamics and be well-adapted to the expanding cosmology in the considered case of de Sitter spacetime.
@@ Acknowledgments
A.T.-C. is grateful to Antonio Vicente-Becerril and Bruno Falcone for constructive discussions on numerics and to the Gravity Laboratory Group of the University of Nottingham for the interesting discussions on the subject. We thank an anonymous referee for helpful presentational suggestions. G.A.M.M. acknowledges financial support from Project No. PID2020-118159GB-C41 and Project No. PID2023-149018NB-C41, both funded by MCIN/AEI/10.13039/501100011033/, and Project No. 2024AEP005 from the Spanish National Research Council (CSIC). The work of J.L. was supported by the United Kingdom Research and Innovation Science and Technology Facilities Council [grant numbers ST/S002227/1, ST/T006900/1 and ST/Y004523/1]. For the purpose of open access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising.